OpenStudy (anonymous):

how do I write a polynomial function when you are given the zeros of that function???

OpenStudy (anonymous):

Write the polynomials having the following zeros: 1) -1, 1, 6 If the polynomial has these zeros, then you can write: x = -1, so x+1 = 0 x = 1, so x-1 = 0 x = 6, so x-6 = 0

OpenStudy (johnweldon1993):

Well you know that when in factored form...a polynomial looks something like (x - something)(x + something) etc... Well if you are given the zeros...you merely see what would make that function = 0 given them Example...if you know the zeros of a function are 3 and -5 (x - 3)(x + 5) why does this work? because when you plug in x = 3 or x = 5 ...that equation = 0

OpenStudy (anonymous):

so i would make those equal to 0? confused...

OpenStudy (zzr0ck3r):

Notice that allot of the work we do with polynomials is a direct result of what we call the zero product rule if ab=0 then a= 0, b=0, or they both equal 0. this is not the case for ab=1 (there are infinite solutions a=1,b=1 or a=1/2 b = 2...... so if we can get from the form of ax^2+bx+c (not there are three terms) into something like this (x-c)(x+d), this is only one term, then we can solve for (x-c)(x+d)=0

OpenStudy (mertsj):

Supposing you have this equation: (x+2)(x-3)=0 Could you solve that?

OpenStudy (anonymous):

oh my lord..... its x^2 -3x+2x-6 = 0???????????????

OpenStudy (mertsj):

x+2=0 or x-3=0 x=-2 or x=3

OpenStudy (mertsj):

Do you understand that?

OpenStudy (mertsj):

It's a simple yes or no question.

OpenStudy (anonymous):

nvm i got it

OpenStudy (mertsj):

Could you do that backwards? What if I said x = -2 and x = 3. What factors produced those answers. Could you tell me?

OpenStudy (anonymous):

x+2 =0 and x-3=0

OpenStudy (mertsj):

Perfect!! So your problem gives you the answers and asks you to write the factors. What answers did it give you?

OpenStudy (anonymous):

2 and -3...

OpenStudy (mertsj):

So, what factors did those answers come from?

OpenStudy (anonymous):

-2 and 3? im getin so confused

OpenStudy (mertsj):

The factors would be x-2 and x+3. The equation would be (x-2)(x+3)=0

OpenStudy (mertsj):

I think part of your problem is that you don't really know what is meant by factors, roots, and the zero factor property.

OpenStudy (anonymous):

im just guessing that i neeed to make em equal to 0 then multiply the expressions and keep it equaled to 0 but thanks for the help?

OpenStudy (mertsj):

Yes. Just as I wrote: (x-2)(x+3)=0 x^2+x-6=0 That is the polynomial equation with the given roots.